The length of a rectangular swimming pool is 2 feet more than twice its width. The perimeter of the pool is 94 feet. What are the length and width of the pool?
Answer by nyc_function(2741)

Let length = 2x + 2
Let width = x
Let P = 94
P = 2L + 2W
94 = 2(2x + 2) + 2x
94 = 4x + 4 + 2x
94 = 6x + 4
94 - 4 = 6x
90 = 6x
90/6 = x
15 = x
The width is 15 feet.
The length is 17 feet.
Done!